Consider a buffer composed of the weak acid acetic acid (CH3COOH) and the conjugate base sodium acetate (NaCH3COO). Which pair of concentrations results in the most effective buffer (i.e. has the highest buffer capacity)?

A. 0.10 M CH3COOH; 0.10 M NaCH3COO

B. 0.90 M CH3COOH; 0.10 M NaCH3COO

C. 0.10 M CH3COOH; 0.90 M NaCH3COO

D. 0.50 M CH3COOH; 0.50 M NaCH3COO
Answer #1

Henderson equation pH = pKa + log [Base / Acid]

for most effective buffer pH = pKa

to get pH = pKa , log [Base/Acid ] = or [Base / Acid ] =

for option D. 0.50 M CH3COOH; 0.50 M NaCH3COO

[NaCH3COO] / [CH3COOH ] = 0.5/ 0.5 = 1

so D will have greater buffer capacity

Answer : D. 0.50 M CH3COOH; 0.50 M NaCH3COO